OpenStudy (needhelpp101):

EEvaluate the function for the given value of x. f(x) = 9x - 6, x = 0 A. f(0) =6 B. f(0) = -9 C. f(0) =-6 D. f(0) = 9

OpenStudy (anonymous):

9(0)-6=0-6=-6

OpenStudy (anonymous):

got it?

OpenStudy (needhelpp101):

no lol

OpenStudy (needhelpp101):

yeah ty :)